(D^2-7DD'+6D'^2)Z=0
1
Expert's answer
2021-09-29T08:45:26-0400

Given "(D^2-7DD'+6D'^2)z = 0"


Let D' = 1, then auxiliary equation will be,

"m^2-7m+6 = 0 \implies (m-1)(m-6) = 0"

"m = 1,6"


Then general solution of the equation will be,

"z = f(y+x)+g(y+6x)"